OpenStudy (anonymous):

A population of bacteria is growing according to the function below, where t is in hours. How many hours will it take for the population to grow to 12,000 bacteria? Round your answer to the nearest integer. Do not include units in your answer.

OpenStudy (campbell_st):

do you have the function...?

OpenStudy (whpalmer4):

@campbell_st where would be the sport in that?! :-)

OpenStudy (campbell_st):

I know... I could come up with a nice exponential model... but the value of the constant is needed...

OpenStudy (anonymous):

Yeah I have it.. sorry B(t) = 50 e2t

terenzreignz (terenzreignz):

\[\Large B(t) = 50e^{2t}\] ...models how many bacteria after t hours, right? So, when the number of bacteria is 12000, then... \[\Large 12000 = 50e^{2t}\] Can you solve for t?

OpenStudy (anonymous):

240/e^2? I don't know how to reduce it much more than that..

terenzreignz (terenzreignz):

Oh... no... I see you divided both sides by 50... everything's fine to that point... \[\Large 240 = e^{2t}\] However, at this point, one does not simply separate the base from exponents... you need the help of the 'natural logarithm' function :)

OpenStudy (whpalmer4):

No, you have to take the natural log of both sides, giving you \[\ln 240=2t\]\[\frac{\ln 240}{2} = t\]and then break out the calculator or your memory of various values of \(\ln\)

terenzreignz (terenzreignz):

Make use of the fact that \[\Large \ln e^{x}=x\]

OpenStudy (anonymous):

I'm sorry, I really have no idea what it means :/ I take an online course that gives you an answer, but not a way to actually figure out the question..

terenzreignz (terenzreignz):

Right. Using this fact... \[\Large \color{blue}\ln \color{red}e^{x} = x\](ln and the exponential function cancel each other out...) And \[\Large 240 = e^{2t}\] You can take the ln of both sides and they'd still be equal... \[\Large \ln 240 = \ln e^{2t}\] Using the property above, what becomes of the right-side?

OpenStudy (anonymous):

I have no idea.. I'm supposed to leave tomorrow and i'm on the final question of my final exam, I could really just use the help :/

terenzreignz (terenzreignz):

See the analogue... \[\Large \ln e^{\color{red}x}=\color{red}x\]\[\Large \ln e^{\color{red}{2t}}=\color{green}?\]

OpenStudy (anonymous):

so.. In e^2t = In240. In e^x = x = In e^2t Which makes the whole equation come out to 240?

OpenStudy (anonymous):

or would 2t = 240?

terenzreignz (terenzreignz):

Slowly... I just pointed out that the ln of e raised to something... is that something... \[\Large \ln e^{x}=x\] Look~ the ln and the e just disappear, leaving whatever it was that e was raised to.

terenzreignz (terenzreignz):

On the other hand, the ln does not 'disappear' in ln 240 because it's not \[\Large \ln e^{240}\]

OpenStudy (anonymous):

Okay, In e^(240) = 240?

terenzreignz (terenzreignz):

Yes, that is true, but that is irrelevant in this question. However, using that idea, what is \[\Large \ln e^{2t}=\color{red}?\]

OpenStudy (anonymous):

Okay, so In e^(2t) = In e^240)?

OpenStudy (anonymous):

which would make the question equal to 120?

terenzreignz (terenzreignz):

There is no \[\Large \ln e^{240}\] involved. I just put it up for the sake of illustrating that... \[\Large \ln 240 \ne 240\]\[\Large \ln e^{240}= 240\]

OpenStudy (anonymous):

Oh okay, one moment then..

OpenStudy (anonymous):

2.7403?

OpenStudy (anonymous):

I did 240 = e^(2t) on a calculator, and that's what it gave me :3

terenzreignz (terenzreignz):

But that's cheating -.- Well, anyway, given that this can't be solved without a calculator, it's correct.

OpenStudy (whpalmer4):

Well, I wouldn't go so far as to say it can't be solved without a calculator. If you've memorized the values of \(\ln 4, \ln 6, \ln 10\), add them up and divide by 2. I find it rather useful to know the first dozen logs and square roots...